OpenStudy (anonymous):

Order of operations problem Is 8 - 2 + 6 = 0 (going left to right) Or is 8 - 2 + 6 = 12 (doing addition first, then subtraction) If you follow the exact order of the letters in Pemdas, you would do the addition first and the answer would be 0. However, if you use convention, you go left to right. Which is it?

OpenStudy (anonymous):

Either you way go it equals 12, you usually work from left to right

OpenStudy (anonymous):

unless pemdas requires the right side first

OpenStudy (anonymous):

This is what I did. 2+6=8 8-8=0

OpenStudy (anonymous):

that is where you messed up it is -2+6=4 and that +8=12

OpenStudy (anonymous):

Let's restate this...is it (8-2) + 6 = 12 or 8 - (2 + 6) = 0 I am helping a 6th grader and there are no parenthesis. I added them.

OpenStudy (anonymous):

Wait so what do you want to know?

OpenStudy (anonymous):

when adding and subtracting problems you can do either or first and in this situation either way you work it the answer is 12 8-2+6=12

OpenStudy (anonymous):

You keep overlooking the fact that the 2 is negative

OpenStudy (anonymous):

Even the teacher added 2 + 6 first, then subtracted the 8. If the 2 is negative, the answer will be 12. If you group the 8 & 2 together, the answer is 0

OpenStudy (anonymous):

I do not think you are allowed to separate the 2 from the - sign, but if the teacher did it then you would do addition first and then subtraction...so by that logic the answer is 0

OpenStudy (anonymous):

The negative 2 makes sense to me but how to explain it to an 11 year old?

OpenStudy (anonymous):

Ok, I was just about to say that; when you separated it 8-(2+6) you would distribute the - all the through. My bad, did not mean to confuse to you and my self .:)

OpenStudy (anonymous):

The negative 2 makes sense to me. And what is the correct way to do PEMDAS when you aren't given the paranthesis? Addition, then subtraction. Or left to right?

OpenStudy (anonymous):

addition then subtraction